OpenStudy (anonymous):

Find the area of the circle in simplest polynomial form. A = 3.14r2 (the radius of the circle is 2x+1) A. 6.28x2 + 3.14 B. 12.56x2 + 3.14> C. 6.28x2 + 3.14x + 3.14 D. 12.56x2 + 12.56x + 3.14 please help

OpenStudy (mertsj):

So I guess the trick here is to square 2x+1 which is the radius. Can you do that?

OpenStudy (anonymous):

A = 3.14r^2 since r = 2x+1 A = 3.14(2x+1)^2 Now all you have to do is simplify this expression. A=3.14(2x+1)(2x+1) A=3.14(4x^2 + 2x + 2x + 1) A=3.14(4x^2 +4x+1) A=12.56x^2 + 12.56x + 3.14